OpenStudy (anonymous):

Which of the following equations would graph a line parallel to 3y=2x-x+5?

OpenStudy (anonymous):

You didn't list the equations. D:

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

3y = x +1 y = x + 1 y = 2/3x + 1 3/2y = x + 1

OpenStudy (anonymous):

Is that "2x-x+5" a typo or is that the actual problem?

OpenStudy (anonymous):

Actual question

OpenStudy (anonymous):

Ok, well, simplify it, and then just find the line with the same slope as that one.

OpenStudy (anonymous):

\[3y = 2x - x + 5\]\[3y = x + 5\]\[y = \frac{x}{3} + \frac{5}{3}\]

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

Basically!

OpenStudy (anonymous):

THanks